Circular No. 6318

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) H. Matthews, Joint Astronomy Centre, Hilo, and Herzberg Institute

of Astrophysics, National Research Council of Canada; and M. Senay and D. Jewitt, University of Hawaii, report the detection of HCN (4-3) emission from comet C/1996 B2 at the James Clerk Maxwell Telescope on Mauna Kea: “Observations on Feb. 10.7 UT show the line at −58.05 ± 0.05 km/s, with peak antenna temperature 0.33 K and area 0.40 ± 0.03 K km/s. The geocentric radial velocity of the comet was −58.01 km/s. The line width (1.1 ± 0.1 km/s) indicates expansion of the coma gases at 550 m/s. Assuming a kinetic temperature of 15 K (30 K), we estimate HCN production rates log Q(HCN) = 26.26 (25.93). This is comparable to the HCN production rate in 1P/Halley at 1.65 AU inbound (Schloerb et al. 1987, A.Ap. 187, 475). The ratio of the HCN to OH production rates (cf. IAUC 6311) is ∼ 0.004. Within the uncertainties due to modelling and to possible temporal variability, this ratio is consistent with the value measured in 1P/Halley at perihelion.”

R. West, European Southern Observatory (ESO), writes that a 15-min CCD exposure obtained by F. Patat on Feb. 16.35 UT with the Danish 1.54-m telescope (+ DFOSC + R filter) shows two broad tails that can be followed from ∼ 3′ from the point of maximum light out to 10′ at the CCD frame’s border. The stronger of the two tails has comparatively straight, sharp edges at p.a. 296^{o} and 286^{o} (i.e., in the general antisun direction) and shows signs of internal structure (‘streamers’); this is probably an ion tail (despite the filter). The other tail is broader, with rather diffuse structure and undefined edges near p.a. 275^{o}. Spectra obtained by Patat on Feb. 16.4 (same telescope) and by E. Cappellaro on Feb. 19.4 (ESO 1.52-m telescope + Boller & Chivens spectrograph) are very similar to that obtained at La Silla on Feb. 8.2 (IAUC 6306), except that on Feb. 19.4 the CN (1,0) band at 917.0 nm is visible.”

GRO J1744–28 M. C. Miller, University of Chicago, reports that the right ascension

of ‘star a’ as given on IAUC 6315 (line 23) should read α = 17^{h}44^{m}36^{s}.72. R. Rutledge and R. Hunstead alerted Vanden Berk et al. that ‘star a’ lies closest to ‘star C’.

1996 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green